# What could go wrong? Discovering and describing failure modes in computer vision

### v1

@@ Abstract
Deep learning models are effective, yet brittle. Even carefully trained, their behavior tends to be hard to predict when confronted with out-of-distribution samples. In this work, our goal is to propose a simple yet effective solution to predict and describe via natural language potential failure modes of computer vision models. Given a pretrained model and a set of samples, our aim is to find sentences that accurately describe the visual conditions in which the model underperforms. In order to study this important topic and foster future research on it, we formalize the problem of Language-Based Error Explainability (LBEE) and propose a set of metrics to evaluate and compare different methods for this task. We propose solutions that operate in a joint vision-and-language embedding space, and can characterize through language descriptions model failures caused, e.g., by objects unseen during training or adverse visual conditions. We experiment with different tasks, such as classification under the presence of dataset bias and semantic segmentation in unseen environments, and show that the proposed methodology isolates nontrivial sentences associated with specific error causes. We hope our work will help practitioners better understand the behavior of models, increasing their overall safety and interpretability.
@@ 1 Introduction
In short, the problem at hand is the following: Given a set of images, denoted by $\mathcal{X}$, a task-specific pre-trained model $\mathcal{M}_{\theta}$, and a large sentence set ${\mathcal{S}}$, our desired output is a subset of sentences ${\mathcal{S}}^{*}\subset{\mathcal{S}}$ that describes in plain text the samples from $\mathcal{X}$ on which the model $\mathcal{M}_{\theta}$ under-performs. To tackle this problem, we propose a simple, yet powerful family of approaches that summarize the visual conditions deemed challenging for a given model via natural language. Our methods rely on uncertainty estimation to distinguish between samples for which the model is more likely to perform well or poorly, and on clustering in a joint vision-and-language embedding space to group together challenging samples associated with specific visual conditions. Operating in this joint representation space allows us to associate sentences to each identified cluster of samples (see illustration in Fig. 1). In order to isolate sentences that are unique to each cluster composed of under-performing samples, we propose different ways to contrast descriptions of those clusters on which the model struggles with those on which it performs performs well (denoted as hard and easy clusters, respectively).
Figure 1: Overview. Provided with a pretrained model ${\mathcal{M}}_{\theta}$, a target image set ${\mathcal{X}}$, and a set of sentences ${\mathcal{S}}$, the family of solutions we propose for LBEE follows the following steps. In Step 1, images from the target set ${\mathcal{X}}$ are split into easy and hard samples, based on the model’s confidence. In Step 2, samples are embedded using Open-CLIP and the visual embeddings of each set are clustered independently. In Step 3, for each hard and easy cluster we compute the cosine similarities between the textual embeddings of candidate sentences and the cluster prototypes. We also assign to each hard prototype the closest easy prototype. Step 4 performs sentence selection for that hard cluster, based on these sentence-prototype similarities. Step 5 aggregates the cluster-specific sentence sets to produce the output.
@@
While the need to provide the sentence set ${\mathcal{S}}$ is a practical limitation of our methods, it plays a crucial role as it allows 1) defining a more generic ground-truth (GT) for evaluation that goes beyond human-in-the-loop qualitative assessment, 2) finding more subtle explanations than clear test set’s biases (e.g.“waterbirds in land” [26], “blonde males” in CelebA [28] or human-implanted biases in NICO++{++}+ +  [9, 51]) 3) performing a fair and quantitative comparisons of different methods, 4) avoiding the need to involve the assessment with a large language model (as e.g.in  [8]).
In summary, the contributions of our paper are: i) formalizing the LBEE problem (Sec. 3); ii) proposing a family of methods to tackle it (Sec. 4); iii) introducing different metrics to evaluate performance; (Sec. 5) iv) showing the effectiveness of our proposed solutions through extensive experiments, focusing on both classification and segmentation segmentation tasks (Sec. 6).
@@ 2 Related work
In particular, Eyuboglu et al. [9, 51] propose to learn a mixture model to partition the data according to classification errors related to spurious correlations. Jain et al. [20] characterize model failures as directions in a latent space found with SVMs. Kim et al. [23] identify biases by analyzing captions from misclassified images with a Vision-Language Model (VLM) and ranks the keywords from the captions using CLIP [32]. Wiles et al. [49] use conditional text-to-image diffusion models to generate synthetic images that are grouped based on how the model misclassifies them. Rezaei et al. [34] extract human-understandable concepts (tags) with VLMs and examine the model’s behavior conditioning on the presence or absence of the combination of those tags. Also Metzen et al. [30] rely on text-to-image models: by encoding specific information in the prompt, they synthesize images from the combinatorially large set of data sub-populations, and run tests to find out on which ones the model under-performs. Finally, a recent work proposes methods to describe differences between two image sets [8]. While this method could be applied to LBEE, the proposed evaluation based on GPT-4 can only assess whether the difference between the sets is correct but not whether this difference is related or not to model failure.
In contrast with the above works, our method only require encoding the images and the textual explanations in a shared visual-textual space such as CLIP, where we rely on simple arithmetic operations among clusters to characterize them via natural language. Our solutions can be seamlessly applied to arbitrary image task—porting this line of work to semantic segmentation for the first time. Differently from previous work on this topic, we also provide a rigorous formalization of the LBEE problem, with ad hoc metrics.
@@
Another body of work tackles interpretability by providing short text descriptions for the reasons a model has provided some output [13, 22, 15, 38]. Such methods are related to our work, in that we are also interested in generating language descriptions for model behaviors; yet, we are interested in failure modes, not in the general rationale behind a prediction.
@@ 3 Language-Based Error Explainability
Letting $\omega^{s_{n}}_{\theta}$ be the average performance of the model on images for which the sentence $s_{n}$ is relevant, we determine that such sentence is describing a failure mode if the difference between the global average $\omega^{\text{avg}}_{\theta}$ (average over all images in ${\mathcal{X}}$) and $\omega^{s_{n}}_{\theta}$ is larger than a threshold $\beta$, with $\beta$ being a predefined margin. The desired output ${\mathcal{S}}^{*}_{\beta}$—used to evaluate the performance of the methods— is the collection of all of these sentences
@@
Relation with prior art. To conclude this section, we position the formulation introduced above with the one tackled by related works [9, 51, 20].
One core difference is that these methods limit their problem formulation to the classification task, while we do not make any assumption related to the nature of the task at hand. Furthermore, in [20] they analyze errors per class, and in [9, 51] assume access to ground truth or prior knowledge about the error types to look for (e.g., bias-conflicting sub-populations or context). In contrast to the above works, we do not use any knowledge about errors prior to evaluation and neither assume—in general—access to class information. Finally, while in prior works the language-based descriptions are part of the proposed methodology, they are not part of the problem formulation and they are not quantitatively evaluated. Instead, we treat the language-based description of errors made by computer vision models as the problem itself, providing means to evaluate them.
@@ 4 A family of approaches to solve LBEE
We propose a family of simple and efficient approaches to tackle the LBEE problem. In contrast with prior art [9, 51, 20], designed for classification and often requiring access to ground-truth or privileged information, our proposed methods are unsupervised and task agnostic. Our solutions rely on two key elements: i) a joint vision-and-language space $\bm{\mathcal{F}}$ , for which we use Open-CLIP [17], and ii) the sentence set ${\mathcal{S}}$ introduced in the previous section, assumed to be large enough to describe images from the target set and to contain potentially relevant reasons for model failure (see Appendix C).
@@
Step 1: Splitting the target set into easy and hard subsets. Given the model ${\mathcal{M}}_{\theta}$ and a confidence measure $\varphi_{\theta}$, we define two thresholds $t_{\varphi}^{h}$ and $t_{\varphi}^{e}$ and split the target set ${\mathcal{X}}$ into three sets: an easy one ${\mathcal{X}}^{e}=\{{\bf x}_{i}\in{\mathcal{X}}|\varphi_{\theta}({\bf x}_{i})>t_{\varphi}^{e}\}$, a hard one ${\mathcal{X}}^{h}=\{{\bf x}_{i}\in{\mathcal{X}}|\varphi_{\theta}({\bf x}_{i})<t_{\varphi}^{h}\}$, and, if $t_{\varphi}^{e}>t_{\varphi}^{h}$, a neutral set with the other images. If ground-truth for the target set is available, we can use the model’s performance to split the data; if not, we can rely on any unsupervised confidence measure.
@@
SetDiff: Sentence set differences. Given the pair of clusters (${\bf c}^{h}_{i}$,${\bf c}^{e}_{j}$), we first select relevant sentences for both—${\mathcal{S}}_{i}^{h}$ and ${\mathcal{S}}_{j}^{e}$—and then remove from ${\mathcal{S}}_{i}^{h}$ the sentences that are present in ${\mathcal{S}}^{e}_{j}$, yielding ${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{h}_{i}\setminus{\mathcal{S}}_{j}^{e}$. Since this approach does not guarantee that $\lvert{\mathcal{R}}_{i}^{h}\rvert=K$, we fill ${\mathcal{S}}_{j}^{e}$ with sentences corresponding to values in ${\bf v}^{e}_{j}$ above a threshold [^2] We set $\tau=0.25$ based on the observation that the average similarity between our sentence sets and image sets is around 0.17 due to the well-known modality gap [25]. $\tau$ and build ${\mathcal{R}}^{h}_{i}$ with the $K$ most similar sentences to ${\bf c}^{h}_{i}$ that are not in ${\mathcal{S}}^{e}_{j}$.
@@
FPDiff: Filtered PDiff. A drawback of PDiff is that it does not guarantee that the selected sentences accurately describe the hard cluster. A remedy to this is to only retain sentences that have high similarity to the prototype ${\bf c}^{h}_{i}$, namely sentences in ${\mathcal{S}}^{d}_{i}$ that are also in ${\mathcal{S}}^{h}_{i}$. We define ${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{h}_{i}\cap{\mathcal{S}}_{i}^{d}$. Also here, to guarantee that $\lvert{\mathcal{R}}_{i}^{h}\rvert=K$, we fill ${\mathcal{S}}_{i}^{h}$ with sentences corresponding to values in ${\bf v}^{h}_{i}$ above a threshold $\tau$ and build ${\mathcal{R}}^{h}_{i}$ with the top $K$ sentences having the highest values in ${\bf d}^{h}_{i}$ that are also in ${\mathcal{S}}^{h}_{i}$.
TopS. We also add a baseline called TopS , describing the hard clusters without any contrasting, by selecting the top $K$ ranked sentences from ${\mathcal{S}}^{h}_{i}$ (${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{h}_{i}$).
@@
Relations with prior art. As mentioned in Sec. 3, previous work related to ours [9, 51] describes a problem formulation that does not perfectly align with ours. Yet, their approaches proposed to assign explanations to predicted error modes can be related to some of the methods we propose above. DOMINO [9], which associates sentences based on average CLIP embedding from which the average class representative is extracted, is closely related to PDiff : when we analyze the errors per class (see Sec. 6), this version of our method can be interpreted as a variant of DOMINO in which the class prototype is replaced with the closest easy prototype. FACTS [51] relies on a captioning tool [31] to assign a single tag to each partition without contrasting, hence, it is mostly related to TopS.
@@ 5 Evaluation metrics for LBEE
Our main goal is retrieving sentences $s_{n}$ that point to reasons for model failure, i.e. $s_{n}\in{\mathcal{S}}^{*}_{\beta}$. This means that the hardness score of the sentence $\omega_{\theta}^{s_{n}}$—formally defined in Eq. 8, Sec. C.2— satisfies the condition $\omega_{\theta}^{s_{n}}<\omega_{\theta}^{avg}-\beta$. The lower this value is, the more the sentence is correlated with model failure. We define the Hardness Ratio (HR) for a given cluster as the ratio of sentences among the retrieved ones in ${\mathcal{R}}_{i}^{h}$ (see Sec. 4) for which the above condition holds, and we define the AHR as the average across all clusters
$\text{HR}_{i}=\frac{\lvert{\mathcal{R}}_{i}^{h}\cap{\mathcal{S}}^{*}_{\beta}\rvert}{\lvert{\mathcal{R}}_{i}^{h}\rvert}\,\,\,\textrm{and}\,\,\,\text{AHR}=\frac{1}{|C^{h}|}\sum_{i=1}^{|C^{h}|}\text{HR}_{i}$ (3)
ii) Average Coverage Ratio (ACR). Next, we introduce a metric that assess the coverage of the sentences associated with the hard clusters, i.e. the ratio of images in the cluster for which the retrieved sentence holds. Intuitively this shows how well a sentence $s_{n}\in{\mathcal{R}}^{h}_{i}$ characterizes a hard cluster ${\bf c}^{h}_{i}$. $\text{CR}_{i}$ is the mean of these ratios taken over the sentences retained in ${\mathcal{R}}^{h}_{i}$. Formally,
@@
where $\Gamma(.,.)$ is a binary operator that takes in input an image and a sentence, and outputs 1 if the sentence $s_{n}$ is relevant for the image ${\bf x}_{m}$ and zero otherwise (see Sec. C.2 for details). To obtain the final ACR value, we simply average $\text{CR}_{i}$ across all clusters, namely
$\text{ACR}=\frac{1}{|C^{h}|}\sum_{i=1}^{|C^{h}|}\text{CR}_{i}.$
@@
$\text{TPR}=\frac{\lvert{\mathcal{R}}_{\mathcal{S}}\cap{\mathcal{S}}^{*}_{\beta}\rvert}{\lvert{\mathcal{S}}^{*}_{\beta}\rvert}\,\,\,\text{and}\,\,\,\text{JI}=\frac{\lvert{\mathcal{R}}_{\mathcal{S}}\cap{\mathcal{S}}^{*}_{\beta}\rvert}{\lvert{\mathcal{R}}_{\mathcal{S}}\cup{\mathcal{S}}^{*}_{\beta}\rvert}.$ (5)

### v2

@@ Abstract
Deep learning models are effective, yet brittle. Even if carefully trained, their behavior tends to be hard to predict when confronted with out-of-distribution samples. In this work, our goal is to propose a simple yet effective solution to predict and describe via natural language potential failure modes of computer vision models. Given a pretrained model and a set of samples, our aim is to find sentences that accurately describe the visual conditions in which the model underperforms. In order to study this important topic and foster future research on it, we formalize the problem of Language-Based Error Explainability (LBEE) and propose a set of metrics to evaluate and compare different methods for this task. We propose solutions that operate in a joint vision-and-language embedding space, and can characterize through language descriptions model failures caused, e.g., by objects unseen during training or adverse visual conditions. We experiment with different tasks, such as classification under the presence of dataset bias and semantic segmentation in unseen environments, and show that the proposed methodology isolates nontrivial sentences associated with specific error causes. We hope our work will help practitioners better understand the behavior of models, increasing their overall safety and interpretability.
@@ 1 Introduction
In short, the problem at hand is the following: Given a set of images, denoted by $\mathcal{X}$, a task-specific pre-trained model $\mathcal{M}_{\theta}$, and a large sentence set ${\mathcal{S}}$, our desired output is a subset of sentences ${\mathcal{S}}^{*}\subset{\mathcal{S}}$ that describes in plain text the samples from $\mathcal{X}$ on which the model $\mathcal{M}_{\theta}$ underperforms. To tackle this problem, we propose a simple, yet powerful family of approaches that summarize the visual conditions deemed challenging for a given model via natural language. Our methods rely on uncertainty estimation to distinguish between samples for which the model is more likely to perform well or poorly, and on clustering in a joint vision-and-language embedding space to group together challenging samples associated with specific visual conditions. Operating in this joint representation space allows us to associate sentences to each identified cluster of samples (see illustration in Fig. 1). In order to isolate sentences that are unique to each cluster composed of under-performing samples, we propose different ways to contrast descriptions of the clusters on which the model struggles with the ones on which it performs performs well (denoted as hard and easy clusters, respectively).
Figure 1: Overview. Provided with a pretrained model ${\mathcal{M}}_{\theta}$, a target image set ${\mathcal{X}}$, and a set of sentences ${\mathcal{S}}$, the family of solutions we propose for LBEE follows the following steps. In Step 1, images from the target set ${\mathcal{X}}$ are split into easy and hard samples, based on the model’s confidence. In Step 2, samples are embedded using Open-CLIP and the visual embeddings of each set are clustered (independently). In Step 3, for each hard and easy cluster we compute the cosine similarities between the textual embeddings of candidate sentences and the cluster prototypes. We also assign the closest easy prototype to each hard prototype. Step 4 performs sentence selection for that hard cluster, based on these sentence-prototype similarities. Step 5 aggregates the cluster-specific sentence sets to produce the output.
@@
While the need to provide the sentence set ${\mathcal{S}}$ is a practical limitation of our methods, it plays a crucial role as it allows 1) defining a more generic ground-truth (GT) for evaluation that goes beyond human-in-the-loop qualitative assessment, 2) finding more subtle explanations than clear test set’s biases (e.g.“waterbirds in land” [26], “blonde males” in CelebA [28] or human-implanted biases in NICO++{++}+ +  [9, 51]), 3) performing a fair and quantitative comparisons of different methods, and 4) avoiding the need to involve the assessment with a large language model (e.g., as in  [8]).
In summary, the contributions of our paper are: i) formalizing the LBEE problem (Sec. 3); ii) proposing a family of methods to tackle it (Sec. 4); iii) introducing different metrics to evaluate performance (Sec. 5) iv); showing the effectiveness of our proposed solutions through extensive experiments, focusing on both classification and semantic segmentation tasks (Sec. 6).
@@ 2 Related work
In particular, Eyuboglu et al. [9, 51] propose to learn a mixture model to partition the data according to classification errors related to spurious correlations. Jain et al. [20] characterize model failures as directions in a latent space learned by SVMs. Kim et al. [23] identify biases by analyzing captions from misclassified images with a Vision-Language Model (VLM) and rank the keywords from the captions using CLIP [32]. Wiles et al. [49] use conditional text-to-image diffusion models to generate synthetic images that are grouped based on how the model misclassifies them. Rezaei et al. [34] extract human-understandable concepts (tags) with VLMs and examine the model’s behavior conditioning on the presence or absence of the combination of those tags. Also Metzen et al. [30] rely on text-to-image models: by encoding specific information in the prompt, they synthesize images from the combinatorially large set of data sub-populations, and run tests to find out on which ones the model underperforms. Finally, a recent work proposes methods to describe differences between two image sets [8]. While this method could be applied to LBEE, the proposed evaluation based on GPT-4 can only assess whether the difference between the sets is correct but not whether this difference is related or not to model failure.
In contrast with the above works, our method only require encoding the images and the textual explanations in a shared visual-textual space such as CLIP, where we rely on simple arithmetic operations among clusters to characterize them via natural language. Our solutions seamlessly apply to arbitrary image task, and indeed we port this line of work to semantic segmentation for the first time. Differently from previous work on this topic, we also provide a rigorous formalization of the LBEE problem, with ad hoc metrics.
@@
Another body of work tackles interpretability by providing short text descriptions to describe why a model has produced some output [13, 22, 15, 38]. Such methods are related to our work, in that we are also interested in generating language descriptions for model behaviors; yet, we are interested in failure modes, not in the general rationale behind a prediction.
@@ 3 Language-Based Error Explainability
Letting $\omega^{s_{n}}_{\theta}$ be the average performance of the model on images for which the sentence $s_{n}$ is relevant, we determine that such sentence is describing a failure mode if the difference between the global average $\omega^{\text{avg}}_{\theta}$ (average over all images in ${\mathcal{X}}$) and $\omega^{s_{n}}_{\theta}$ is larger than a threshold $\beta$, namely a predefined margin. The desired output ${\mathcal{S}}^{*}_{\beta}$, used to evaluate the performance of the methods, is the collection of all of these sentences
@@
Relation with prior art. To conclude this section, we position the formulation introduced above with the ones tackled by related works [9, 51, 20].
One core difference is that these methods limit their problem formulation to the classification task, while we do not make any assumption related to the nature of the task at hand. Furthermore, in [20] they analyze errors per class, and in [9, 51] assume access to ground truth or prior knowledge about the error types to look for (e.g., bias-conflicting sub-populations or context). In contrast to the above works, we do not use any knowledge about errors prior to evaluation and neither assume, in general, access to class information. Finally, while in prior works the language-based descriptions are part of the proposed methodology, they are not part of the problem formulation and they are not quantitatively evaluated. Instead, we treat the language-based description of errors made by computer vision models as the problem itself, providing means to evaluate them.
@@ 4 A family of approaches to solve LBEE
We propose a family of simple, unsupervised and task agnostic approaches to tackle the LBEE problem. Our solutions rely on two key elements: i) a joint vision-and-language space $\bm{\mathcal{F}}$ , for which we use Open-CLIP [17], and ii) the sentence set ${\mathcal{S}}$ introduced in the previous section, assumed to be large enough to describe images from the target set and to contain potentially relevant reasons for model failure (see Appendix C).
@@
Step 1: Splitting the target set into easy and hard subsets. Given the model ${\mathcal{M}}_{\theta}$ and a confidence measure $\varphi_{\theta}$, we define two thresholds $t_{\varphi}^{h}$ and $t_{\varphi}^{e}$ and split the target set ${\mathcal{X}}$ into three sets: an easy one ${\mathcal{X}}^{e}=\{{\bf x}_{i}\in{\mathcal{X}}|\varphi_{\theta}({\bf x}_{i})>t_{\varphi}^{e}\}$, a hard one ${\mathcal{X}}^{h}=\{{\bf x}_{i}\in{\mathcal{X}}|\varphi_{\theta}({\bf x}_{i})<t_{\varphi}^{h}\}$, and, if $t_{\varphi}^{e}>t_{\varphi}^{h}$, a neutral set with the remaining images. If ground truth for the target set is available, we can instead use the model’s performance to split the data.
@@
SetDiff: Sentence set differences. Given a pair of clusters (${\bf c}^{h}_{i}$,${\bf c}^{e}_{j}$), we first select relevant sentences for both—${\mathcal{S}}_{i}^{h}$ and ${\mathcal{S}}_{j}^{e}$—and then remove from ${\mathcal{S}}_{i}^{h}$ the sentences that are present in ${\mathcal{S}}^{e}_{j}$, yielding ${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{h}_{i}\setminus{\mathcal{S}}_{j}^{e}$. Since this approach does not guarantee that $\lvert{\mathcal{R}}_{i}^{h}\rvert=K$, we fill ${\mathcal{S}}_{j}^{e}$ with sentences corresponding to values in ${\bf v}^{e}_{j}$ above a threshold [^2] We set $\tau=0.25$ based on the observation that the average similarity between our sentence sets and image sets is around 0.17, due to the well-known modality gap [25]. $\tau$ and build ${\mathcal{R}}^{h}_{i}$ with the $K$ most similar sentences to ${\bf c}^{h}_{i}$ that are not in ${\mathcal{S}}^{e}_{j}$.
@@
FPDiff: Filtered PDiff. A drawback of PDiff is that it does not guarantee that the selected sentences accurately describe the hard cluster. A remedy to this is only retaining sentences that have high similarity to the prototype ${\bf c}^{h}_{i}$, namely sentences in ${\mathcal{S}}^{d}_{i}$ that are also in ${\mathcal{S}}^{h}_{i}$. We define ${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{h}_{i}\cap{\mathcal{S}}_{i}^{d}$. Also here, to guarantee that $\lvert{\mathcal{R}}_{i}^{h}\rvert=K$, we fill ${\mathcal{S}}_{i}^{h}$ with sentences corresponding to values in ${\bf v}^{h}_{i}$ above a threshold $\tau$ and build ${\mathcal{R}}^{h}_{i}$ with the top $K$ sentences having the highest values in ${\bf d}^{h}_{i}$ that are also in ${\mathcal{S}}^{h}_{i}$.
TopS. We also include a baseline called TopS , describing the hard clusters without any contrasting, by selecting the top $K$ ranked sentences from ${\mathcal{S}}^{h}_{i}$ (${\mathcal{R}}^{h}_{i}={\mathcal{S}}^{h}_{i}$).
@@
Relations with prior art. As mentioned in Sec. 3, prior art [9, 51] describes a problem formulation that does not perfectly align with ours. Yet, their approaches proposed to assign explanations to predicted error modes can be related to some of the methods we propose above. DOMINO [9], which associates sentences based on average CLIP embedding from which the average class representative is extracted, is closely related to PDiff : when we analyze the errors per class (see Sec. 6), this version of our method can be interpreted as a variant of DOMINO in which the class prototype is replaced with the closest easy prototype. FACTS [51] relies on a captioning tool [31] to assign a single tag to each partition without contrasting, hence, it is mostly related to TopS.
@@ 5 Evaluation metrics for LBEE
i) Average Hardness Ratio (AHR). Our main goal is retrieving sentences $s_{n}$ that point to reasons for model failure, i.e., $s_{n}\in{\mathcal{S}}^{*}_{\beta}$. This means that the hardness score of the sentence $\omega_{\theta}^{s_{n}}$—formally defined in Eq. 8, Sec. C.2— satisfies the condition $\omega_{\theta}^{s_{n}}<\omega_{\theta}^{avg}-\beta$. The lower this value is, the more the sentence is correlated with model failure.
We define the Hardness Ratio (HR) for a given cluster as the ratio of sentences among the retrieved ones in ${\mathcal{R}}_{i}^{h}$ (see Sec. 4) for which the above condition holds
$\text{HR}_{i}=\frac{\lvert{\mathcal{R}}_{i}^{h}\cap{\mathcal{S}}^{*}_{\beta}\rvert}{\lvert{\mathcal{R}}_{i}^{h}\rvert}\enspace,$ (3)
and we define the AHR as the average across all clusters
$\text{AHR}=\frac{1}{|C^{h}|}\sum_{i=1}^{|C^{h}|}\text{HR}_{i}\enspace.$
ii) Average Coverage Ratio (ACR). Next, we introduce a metric to assess the coverage of the sentences associated with the hard clusters, i.e. the ratio of images in the cluster for which the retrieved sentence holds. Intuitively this shows how well a sentence $s_{n}\in{\mathcal{R}}^{h}_{i}$ characterizes a hard cluster ${\bf c}^{h}_{i}$. $\text{CR}_{i}$ is the mean of these ratios taken over the sentences retained in ${\mathcal{R}}^{h}_{i}$. Formally,
@@
where $\Gamma(.,.)$ is a binary operator that, taking in input an image and a sentence, outputs 1 if the sentence $s_{n}$ is relevant for the image ${\bf x}_{m}$ and zero otherwise (see Sec. C.2 for details). To obtain the final ACR value, we simply average $\text{CR}_{i}$ across all clusters, namely
$\text{ACR}=\frac{1}{|C^{h}|}\sum_{i=1}^{|C^{h}|}\text{CR}_{i}\enspace.$
@@
$\text{TPR}=\frac{\lvert{\mathcal{R}}_{\mathcal{S}}\cap{\mathcal{S}}^{*}_{\beta}\rvert}{\lvert{\mathcal{S}}^{*}_{\beta}\rvert}\,\,\,\text{and}\,\,\,\text{JI}=\frac{\lvert{\mathcal{R}}_{\mathcal{S}}\cap{\mathcal{S}}^{*}_{\beta}\rvert}{\lvert{\mathcal{R}}_{\mathcal{S}}\cup{\mathcal{S}}^{*}_{\beta}\rvert}\enspace.$ (5)